If I subtract from twicemy present age three times my age six years ago, I get my present. What is my age ?
Answer by sachi(548)

let my present age=x
six yrs ago=x-6
now 2x-3(x-6)=x
or 2x-3x+18=x
or 2x=18 or x=9 yrs
ans